OpenStudy (anonymous):

Write a rule for the linear function in the table. x f(x) 1 –7 2 –10 3 –13 4 –16

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} x\qquad &y=f(x)\\ \hline\\ 1&-7\\ 2&-10\implies -7-3\\ 3&-13\implies -10-3\\ 4&-16\implies -13-3 \end{array}\) what do you think?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

-7,-10,-13,-16, it is an A.P. here a=-7,d=-10-(-7)=-10+7=-3 tn=a+(n-1)d tn=-7+(n-1)(-3)=-7-3n+3=-3n-4 Replace n by x and tn by y y=-3x-4,x=1,2,3,...

OpenStudy (anonymous):

correction write f(x) for y